3. You have found three investment choices for a​ one-year deposit: 9.9% APR compounded​ monthly, 9.9% APR compounded​ annually, and 9.4% APR compounded daily. Compute the EAR for each investment choice.​ (Assume that there are 365 days in the​ year.) ​(Note: Be careful not to round any intermediate steps less than six decimal​ places.)
Answer #1

EAR for 9.9% APR compounded​ monthly:

EAR = (1 + APR/n)n - 1

EAR = (1 + 0.099/12)12 - 1

EAR = (1 + 0.00825)12 - 1

EAR = 0.1036 - 1

EAR = 0.1036 or 10.36%

EAR for 9.9% APR compounded​ annually:

EAR = (1 + APR/n)n - 1

EAR = (1 + 0.099/1)1 - 1

EAR = (1 + 0.099)1 - 1

EAR = 0.099 - 1

EAR = 0.099 or 9.9%

EAR for 9.4% APR compounded​ daily:

EAR = (1 + APR/n)n - 1

EAR = (1 + 0.094/365)365 - 1

EAR = (1 + 0.000258)365 - 1

EAR = 1.0985 - 1

EAR = 0.0985 or 9.85%